# Towards Graph Foundation Models: A Study on the Generalization of Positional and Structural Encodings

### v1

@@ 1 Introduction
In recent years, graph neural networks (GNNs) and machine learning on graphs have gained significant attention due to their versatility in processing complex data across various domains, including biology, social networks, chemistry, and recommendation systems. Comprehensive surveys on these advancements are available in Bronstein et al. [1] and Waikhom and Patgiri [2]. GNNs are fundamentally designed to extend traditional deep learning techniques to graph-structured data by aggregating information from local neighborhoods, thereby enabling the extraction of structural and relational features intrinsic to graphs. Despite considerable progress, these methods continue to face several challenges, including issues related to expressivity, such as the ability to distinguish between non-isomorphic graphs [3] and substructure counting [4], the generalization of models across different graph datasets and tasks [5], as well as the connection between generalization and expressivity that has been studied recently [6].
@@
In terms of expressivity, i.e., the ability of GNNs to approximate functions [7], message-passing neural networks (MPNNs) are known to have limitations that can hinder their performance in practical applications [3, 8, 9, 10]. To address these limitations, various strategies have been proposed, with two prominent approaches focusing on augmenting the input node features of a graph with positional and structural encodings (PE and SE, respectively). One such approach involves using random node features (RNF) [11, 12], which have been shown to enhance the expressivity of MPNNs by increasing their ability to distinguish between nodes. However, despite the theoretical benefits, RNF has not consistently improved performance on real-world datasets [4]. To address this inconsistency, recent studies have demonstrated that learned PEs [13] and learned graph normalization techniques [14] can leverage RNF to improve both expressivity and downstream task performance. Another approach to overcoming the expressivity limitations involves incorporating spectral PEs, which are often derived from a partial eigendecomposition of the graph-Laplacian matrix [15, 16, 17, 18, 19]. Additionally, structural encodings based on random walks [20] have also proven to be beneficial. More recently, the graph positional and structural encoder (GPSE) [21] introduced a unified framework for learning from both positional and structural encodings while utilizing RNF. Overall, these methods have shown improvements, particularly on specific graph benchmarks where models are trained from scratch for each task. Notably, GPSE also demonstrated some degree of generalization to new tasks.
@@
Simultaneously with the advancements in GNNs, foundation models (FM) have emerged and revolutionized the fields of natural language processing (NLP) and computer vision (CV). These developments present an exciting opportunity to bring similar transformative advances to machine learning on graphs. In NLP, models like BERT [22] and GPT [23] have demonstrated how pre-training on large datasets can produce models that generalize effectively to a variety of downstream tasks with minimal to no fine-tuning. Inspired by these successes, recent works [24, 25, 26, 27, 28, 29, 30] have begun exploring the creation of graph foundation models (GFM), which could offer similar benefits for graph-based tasks. Such models would not only simplify the training process across multiple tasks but also significantly reduce the computational resources, energy consumption, and data requirements typically associated with task-specific training. This aligns with the broader objective of developing more environmentally sustainable machine learning practices.
@@
Our Contributions. In this paper, we both practically and theoretically assess whether PSEs can function as foundation models or, alternatively, as a building block for future foundation models. We especially focus on the recent GPSE model which we consider the state of the art when it comes to PSEs. We note that, our main contribution is the study of the role of GPSE as a building block within future Graph Foundation Models, rather than proposing a new methodology. We believe that these contributions are useful and important for the graph-learning community. To be precise we:
@@
Our findings indicate that while GPSE and other PSEs do not yet function as standalone GFMs due to their performance being somewhat dependent on the dataset in question, they show potential as valuable components in future GFMs due to their increased robustness in data-scarce settings and their improved generalization. This potential is primarily attributed to GPSE’s strong generalization capabilities, which we identify as a critical factor for enhancing downstream performance (Section 5.3). These results underscore the importance of studying generalization in graph learning, a key area for advancing the development and understanding of GFMs, as also highlighted by Morris et al. [31].
@@ 2 Related Work
Graph foundation models. Inspired by the success of FMs in NLP [23] and CV [32], there has been increasing interest in developing similar models for graph-based tasks. These GFMs aim to generalize across various graph-related tasks by leveraging large-scale pre-training and adaptable architectures. Liu et al. [24] provide a comprehensive analysis of the current state of GFMs. Their study highlights significant challenges in the development of general GFMs, noting that the diversity of graph structures and tasks makes it difficult to create models with broad applicability. Shi et al. [28] emphasize the potential impact of GFMs on a wide range of applications, from social network analysis to molecular modeling. Also, Mao et al. [27] discuss the potential directions for designing effective GFMs, and propose a taxonomy of task-specific GFMs, outlining the key design principles required to achieve generalization across various graph tasks. Xia et al. [29], Liu et al. [30], and Huang et al. [26] propose GFMs by use of graph-tokenization, text-attributed graphs, and in-context learning for graphs. These models have difficulties with balancing task specificity with generalization capabilities or regression tasks. A promising example of a domain-specific GFM is presented in Galkin et al. [25], a GFM tailored for knowledge graph tasks was shown to be successful, excelling in tasks such as knowledge graph completion and link prediction, showcasing how specialized GFMs can achieve state-of-the-art results in specific domains. However, there is still a need to develop GFMs for other domains like biology and chemistry, where PSEs are commonly applied.
@@
On a different note, recently, there has been a series of works that utilize large language models (LLMs) as FMs for graph learning tasks, such as [33, 34] and others. We note that, this direction is very promising in terms of the large availability of pre-trained models, as well as the abundance of textual data used in LLMs, compared with the scarcity of graph data. However, in this paper, we focus on employing GNNs and concepts within, to study whether and how they can be oriented towards GFMs, which can be later combined with LLM approaches.
@@
Positional and structural encodings PSEs are features that provide information on the graph’s structure and node positions. The seminal work in Vaswani et al. [35] introduced the concept of positional encodings in the transformer architecture, demonstrating their effectiveness in capturing the order of sequences. In graph learning, the inclusion of PSEs was shown to be significant for enhancing the performance of models in various tasks, as we now discuss. Methods utilizing graph Laplacian eigenvectors as PEs (LapPE) have been explored by Kreuzer et al. [16] and Rampášek et al. [19]. Kreuzer et al. [16] also investigated the use of electrostatic potential encodings, providing a novel approach to structural representation. Ying et al. [36] proposed using shortest paths as encodings, which have shown promise in capturing graph structures. Shiv and Quirk [37] introduced tree-based encodings, offering a unique perspective on hierarchical data representation. Random walk-based encodings (RWSE) have been explored by Rampášek et al. [19], Dwivedi and Bresson [15] and Dwivedi et al. [38], highlighting their versatility in various applications. The use of heat kernels for SEs has been discussed by Kreuzer et al. [16] and Mialon et al. [39]. Subgraph-based encodings, which have been effective in capturing local graph structures, were the focus of Zhao et al. [40], Bouritsas et al. [41], Chen et al. [42] and Yan et al. [43]. Additionally, Ying et al. [36] considered node degree as a simple yet powerful SEs. Furthermore, a different set of works has shown that RNF [11, 12, 44] are theoretically strong PEs, and they were further utilized in Eliasof et al. [13] for learnable PEs via their propagation. Recently, it was shown in Cantürk et al. [21] that RNF can be combined with the aforementioned PSEs to learn a general encoding, called graph positional and structural encoding (GPSE). While PSEs have shown great promise in improving performance on various tasks, their supposed effects on GFMs have not yet been studied in detail.
@@
Expressivity of GNNs The expressivity of GNNs has been widely studied in recent years [8, 3, 45], highlighting the theoretical limits of GNNs. To improve the expressivity of GNNs, several approaches were proposed and studied, from substructure encodings [41], subgraph GNNs [4, 40, 46], to homomorphism counting [47], to the augmentation of input node features with PSEs such as RNF [48, 49, 12, 11, 44]. In this work, following GPSE that utilizes RNF as part of its architecture, we theoretically discuss the expressive power of GPSE, and study whether RNF are required in practice for downstream tasks when utilizing a framework such as GPSE. A further discussion of expressivity not considered here can be found in Appendix A.
@@ 3 Notations and Background
Color refinement (CR) CR is a well-studied heuristic for the graph isomorphism problem, originally proposed by Leman and Weisfeiler [50]. Intuitively, the algorithm determines if two graphs are non-isomorphic by iteratively coloring or labeling vertices. Formally, let be a labeled graph. In each iteration, , CR computes a vertex coloring , depending on the coloring of the neighbors. That is, in iteration , we set
@@
Message-passing neural networks (MPNNs) Intuitively, MPNNs learn a vectorial representation, i.e., a -dimensional real-valued vector, representing each node in a graph by aggregating information from neighboring vertices. Formally, let be a labeled graph with initial node features . A typical MPNN architecture is obtained by stacking permutation-equivariant parameterized neural layers. Following, Scarselli et al. [51] and Gilmer et al. [52], in each layer, , we define the updated node features as:
for each . The functions and are typically learned. Two examples of common MPNN layers, that will be considered in this work, are the graph isomorphism network (GIN) layer [8] defined as:
where is a multilayer perceptron (MLP) in layer and is a constant of layer differentiating a node from its neighbors. Further, the gated graph convolutional network (GatedGCN) layer [53] is defined as:
@@
Graph positional and structural encoder (GPSE) The GPSE model proposed in Cantürk et al. [21] is the main architecture studied in this work. Given the structure of a graph , namely without any labels, as well as some random features it provides node embeddings It is pre-trained on MolPCBA to predict various PSEs with prediction heads attached. Throughout most of this paper we consider GPSE to be pre-trained and thus fixed. We thus think of GPSE as a PSE that provides node embeddings, and provide additional details in Appendix B.
@@ 4.1 GPSE is a Universal Node Encoder
Since GPSE is meant to be able to compute PSEs for any input graph and thus provide a powerful positional and structural node encoding, it is important to verify that GPSE is indeed a universal node encoder. Because some PSEs are harder—in a WL sense—to compute than CR, GPSE as an MPNN needs to be fully expressive. And indeed, GPSE is a universal node encoder, because the input graph is augmented with RNF, which we now show. First, note that GatedGCN layers can approximate GIN layers by the following theorem, which we prove in Appendix C.
@@
Combining Theorem 1 with Theorem 3 from Xu et al. [8], which states that an MPNN made up of GIN layers can learn to approximate 1-WL, proves that GatedGCN can also simulate 1-WL. Furthermore, Theorem 4 from Franks et al. [44] shows that, assuming sufficiently many layers and sufficient width of the layers within GPSE, any function on the graph structure can be computed. This proves that, indeed, under mild assumptions, GPSE is universal and can learn to embed any PSE.
@@ 4.2 PSEs do not Guarantee Downstream Universality
While GPSE can learn to embed any PSE, this is not sufficient to ensure the universality of the downstream model. In fact, PSEs in general do not provide universality of the downstream model. As shown in Cantürk et al. [21], as well as in Table 1, a GIN downstream network augmented with GPSE is not able to accurately generalize on the EXP, CEXP, CSL, TRI, and TRIX datasets, that are often used as expressivity benchmarks [11, 49]. Let us assume that the pre-trained network in GPSE learned an identity mapping of the input random node features, then the resulting model is theoretically universal with high probability, as shown in Abboud et al. [11] and Sato et al. [12]. This would imply that the issue lies in the pre-training of GPSE. That is, GPSE did not learn to encode PSEs which obtain high accuracy on some graphs that require high expressivity to distinguish.
GPSE+. We empirically verify that by adding relevant “hard” graphs, i.e., graphs that require high expressivity, to the pre-training dataset of GPSE, GPSE is practically more expressive. Specifically, relevant hard graphs for the CSL dataset are 4-regular graphs. Thus we define GPSE+to be the exact same model as GPSE trained on MolPCBA (¿300.000 unique graphs), as well as 1000 random 4-regular graphs, each with 24 nodes. We chose 24 nodes, as this is roughly the average number of nodes for MolPCBA. Indeed, Table 1 shows that, by adding these hard graphs to the pre-training stage, the performance of GPSE on CSL increases significantly. The performance on EXP also increases; however, since the added graphs are not particularly relevant to the EXP dataset, the performance does not change as dramatically as in the case of CSL. We also evaluated on TRI and TRIX, which are 3-regular graphs with triangles as labels, where again GPSE+outperforms GPSE.
@@
GPSE-. While the GPSE+variant considers strengthening the pre-training stage with hard graphs, our experimental results in Section 5 on real-world datasets do not read improved performance. Therefore, to address the question of the importance of expressivity in GPSE, we also replace the input RNF with a constant feature vector, effectively removing the universality offered by RNF. We call this variant GPSE-. It has the same model architecture as GPSE, however, instead of random normally distributed inputs it uses vectors, i.e. . In practice, this implies that the GPSE-model is not capable of computing all PSEs for any input graph. As an example, GPSE-cannot compute cycle counts for the graphs in Fig. 1(a) and Fig. 1(b), since GPSE-cannot tell these two graphs apart. However, since molecules are almost trees [54], GPSE-should be able to still learn the target PSEs for the pre-training dataset, MolPCBA, as well as for the other evaluation datasets like ZINC-12k. Notably, Table 1 shows that GPSE-does not excel on datasets that involve graphs harder than 1-WL. In particular, the performance of GPSE-is roughly in line with a constant model. On TRIX GPSE-outperforms GPSE as well as GPSE+, because these other variants are having trouble generalizing well to the larger 3-regular graphs of TRIX, and actually perform worse than a naive, constant predictor. The evaluations in Section 5 will demonstrate that, indeed, RNF is not necessary for good practical performance either. In practice, at least for current datasets, GPSE, GPSE+, and GPSE-perform comparably.

### v2

@@ 1 Introduction
In recent years, graph neural networks (GNNs) and machine learning on graphs have gained significant attention due to their versatility in processing complex data across various domains, including biology, social networks, chemistry, and recommendation systems. Comprehensive surveys on these advancements are available in Bronstein et al. (2021) and Waikhom and Patgiri (2023). GNNs are fundamentally designed to extend traditional deep learning techniques to graph-structured data by aggregating information from local neighborhoods, thereby enabling the extraction of structural and relational features intrinsic to graphs. Despite considerable progress, these methods continue to face several challenges, including issues related to expressivity, such as the ability to distinguish between non-isomorphic graphs (Morris et al., 2019) and substructure counting (Bevilacqua et al., 2022), the generalization of models across different graph datasets and tasks (Ju et al., 2023), as well as the connection between generalization and expressivity that has been studied recently (Li et al., 2024).
@@
In terms of expressivity, i.e., the ability of GNNs to approximate functions (Huang and Villar, 2021), message-passing neural networks (MPNNs) are known to have limitations that can hinder their performance in practical applications (Morris et al., 2019; Xu et al., 2019; Morris et al., 2023; Zhang et al., 2021). To address these limitations, various strategies have been proposed, with two prominent approaches focusing on augmenting the input node features of a graph with positional and structural encodings (PE and SE, respectively). One such approach involves using random node features (RNF) (Abboud et al., 2021; Sato et al., 2021), which have been shown to enhance the expressivity of MPNNs by increasing their ability to distinguish between nodes. However, despite the theoretical benefits, RNF has not consistently improved performance on real-world datasets (Bevilacqua et al., 2022; Bechler-Speicher et al., 2025). To address this inconsistency, recent studies have demonstrated that learned PEs (Eliasof et al., 2023) and learned graph normalization techniques (Eliasof et al., 2024) can leverage RNF to improve both expressivity and downstream task performance. Another approach to overcoming the expressivity limitations involves incorporating spectral PEs, which are often derived from a partial eigendecomposition of the graph-Laplacian matrix (Dwivedi and Bresson, 2021; Kreuzer et al., 2021; Wang et al., 2022a; Lim et al., 2023; Rampášek et al., 2022). Additionally, structural encodings based on random walks (Dwivedi et al., 2022a) have also proven to be beneficial. More recently, the graph positional and structural encoder (GPSE) (Cantürk et al., 2024) introduced a unified framework for learning from both positional and structural encodings while utilizing RNF. Overall, these methods have shown improvements, particularly on specific graph benchmarks where models are trained from scratch for each task. Notably, GPSE also demonstrated some degree of generalization to new tasks.
@@
Simultaneously with the advancements in GNNs, foundation models (FM) have emerged and revolutionized the fields of natural language processing (NLP) and computer vision (CV). These developments present an exciting opportunity to bring similar transformative advances to machine learning on graphs. In NLP, models like BERT (Devlin et al., 2019) and GPT (Brown et al., 2020) have demonstrated how pre-training on large datasets can produce models that generalize effectively to a variety of downstream tasks with minimal to no fine-tuning. Inspired by these successes, recent works (Liu et al., 2023; Galkin et al., 2024; Huang et al., 2023a; Mao et al., 2024; Shi et al., 2024; Xia et al., 2024; Liu et al., 2024) have begun exploring the creation of graph foundation models (GFM), which could offer similar benefits for graph-based tasks. Such models would not only simplify the training process across multiple tasks but also significantly reduce the computational resources, energy consumption, and data requirements typically associated with task-specific training. This aligns with the broader objective of developing more environmentally sustainable machine learning practices.
@@
Our Contributions. In this paper, we both practically and theoretically assess whether PSEs can function as foundation models or, alternatively, as a building block for future foundation models. We especially focus on the recent GPSE model which we consider the state of the art when it comes to PSEs. We note that, our main contribution is the study of the role of GPSE as a building block within future GFMs, rather than proposing a new methodology. We believe that these contributions are useful and important for the graph-learning community. To be precise we:
@@
Our findings indicate that while GPSE and other PSEs do not yet function as standalone GFMs due to their performance being somewhat dependent on the dataset in question, they show potential as valuable components in future GFMs due to their increased robustness in data-scarce settings and their improved generalization. This potential is primarily attributed to GPSE’s strong generalization capabilities, which we identify as a critical factor for enhancing downstream performance (Section 5.3). These results underscore the importance of studying generalization in graph learning, a key area for advancing the development and understanding of GFMs, as also highlighted by Morris et al. (2024).
@@ 2 Related Work
Graph foundation models. Inspired by the success of FMs in NLP (Brown et al., 2020) and CV (Dosovitskiy et al., 2021), there has been increasing interest in developing similar models for graph-based tasks. These GFMs aim to generalize across various graph-related tasks by leveraging large-scale pre-training and adaptable architectures. Liu et al. (2023) provide a comprehensive analysis of the current state of GFMs. Their study highlights significant challenges in the development of general GFMs, noting that the diversity of graph structures and tasks makes it difficult to create models with broad applicability. Shi et al. (2024) emphasize the potential impact of GFMs on a wide range of applications, from social network analysis to molecular modeling. Also, Mao et al. (2024) discuss the potential directions for designing effective GFMs, and propose a taxonomy of task-specific GFMs, outlining the key design principles required to achieve generalization across various graph tasks. Xia et al. (2024); Liu et al. (2024), and Huang et al. (2023a) propose GFMs by use of graph-tokenization, text-attributed graphs, and in-context learning for graphs. These models have difficulties with balancing task specificity with generalization capabilities or regression tasks. A promising example of a domain-specific GFM is presented in Galkin et al. (2024), a GFM tailored for knowledge graph tasks was shown to be successful, excelling in tasks such as knowledge graph completion and link prediction, showcasing how specialized GFMs can achieve state-of-the-art results in specific domains. However, there is still a need to develop GFMs for other domains like biology and chemistry, where PSEs are commonly applied, as recently discussed in Frasca et al. (2024).
@@
On a different note, recently, there has been a series of works that utilize large language models (LLMs) as FMs for graph learning tasks, such as in Fatemi et al. (2024); Perozzi et al. (2024) and others. We note that, this direction is very promising in terms of the large availability of pre-trained models, as well as the abundance of textual data used in LLMs, compared with the scarcity of graph data. However, in this paper, we focus on employing GNNs and concepts within, to study whether and how they can be oriented towards GFMs, which can be later combined with LLM approaches.
@@
Positional and structural encodings. PSEs are features that provide information on the graph’s structure and node positions. The seminal work in Vaswani et al. (2017) introduced the concept of positional encodings in the transformer architecture, demonstrating their effectiveness in capturing the order of sequences. In graph learning, the inclusion of PSEs was shown to be significant for enhancing the performance of models in various tasks (Srinivasan and Ribeiro, 2020), as we now discuss. Methods utilizing graph Laplacian eigenvectors as PEs (LapPE) have been explored by Kreuzer et al. (2021) and Rampášek et al. (2022). Kreuzer et al. (2021) also investigated the use of electrostatic potential encodings, providing a novel approach to structural representation. Ying et al. (2021) proposed using shortest paths as encodings, which have shown promise in capturing graph structures. Shiv and Quirk (2019) introduced tree-based encodings, offering a unique perspective on hierarchical data representation. Random walk-based encodings (RWSE) have been explored by Rampášek et al. (2022); Dwivedi and Bresson (2021) and Dwivedi et al. (2022b), highlighting their versatility in various applications. The use of heat kernels for SEs has been discussed by Kreuzer et al. (2021) and Mialon et al. (2021). Subgraph-based encodings, which have been effective in capturing local graph structures, were the focus of Zhao et al. (2022); Bouritsas et al. (2022); Chen et al. (2022) and Yan et al. (2023). Additionally, Ying et al. (2021) considered node degree as a simple yet powerful SEs. Furthermore, a different set of works has shown that RNF (Abboud et al., 2021; Sato et al., 2021; Franks et al., 2023) are theoretically strong PEs, and they were further utilized in Eliasof et al. (2023) for learnable PEs via their propagation. Recently, it was shown in Cantürk et al. (2024) that RNF can be combined with the aforementioned PSEs to learn a general encoding, called graph positional and structural encoding (GPSE). While PSEs have shown great promise in improving performance on various tasks, their supposed effects on GFMs has been under-explored, with preliminary experimental results shown in Frasca et al. (2024). In this paper, we offer a detailed study on the generalization ability of learned PSEs, accompanied by a theoretical discussion of their expressiveness.
@@
Expressivity of GNNs The expressivity of GNNs has been widely studied in recent years (Xu et al., 2019; Morris et al., 2019; Maron et al., 2019a), highlighting the theoretical limits of GNNs. To improve the expressivity of GNNs, several approaches were proposed and studied, from substructure encodings (Bouritsas et al., 2022), subgraph GNNs (Bevilacqua et al., 2022; Zhao et al., 2022; Bevilacqua et al., 2024), to homomorphism counting (Zhang et al., 2024), to the augmentation of input node features with PSEs such as RNF (Dasoulas et al., 2020; Murphy et al., 2019; Sato et al., 2021; Abboud et al., 2021; Franks et al., 2023). In this work, following GPSE that utilizes RNF as part of its architecture, we theoretically discuss the expressive power of GPSE, and study whether RNF are required in practice for downstream tasks when utilizing a framework such as GPSE. A further discussion of expressivity not considered here can be found in Appendix A.
@@ 3 Notations and Background
Color refinement (CR) CR is a well-studied heuristic for the graph isomorphism problem, originally proposed by Leman and Weisfeiler (1968). Intuitively, the algorithm determines if two graphs are non-isomorphic by iteratively coloring or labeling vertices. Formally, let be a labeled graph. In each iteration, , CR computes a vertex coloring , depending on the coloring of the neighbors. That is, in iteration , we set
@@
Message-passing neural networks (MPNNs) Intuitively, MPNNs learn a vectorial representation, i.e., a -dimensional real-valued vector, representing each node in a graph by aggregating information from neighboring vertices. Formally, let be a labeled graph with initial node features . A typical MPNN architecture is obtained by stacking permutation-equivariant parameterized neural layers. Following, Scarselli et al. (2008) and Gilmer et al. (2017), in each layer, , we define the updated node features as:
for each . The functions and are typically learned. Two examples of common MPNN layers, that will be considered in this work, are the graph isomorphism network (GIN) layer (Xu et al., 2019) defined as:
where is a multilayer perceptron (MLP) in layer and is a constant of layer differentiating a node from its neighbors. Further, the gated graph convolutional network (GatedGCN) layer (Bresson and Laurent, 2017) is defined as:
@@
3.1 Graph positional and structural encoder (GPSE)
GPSE (Cantürk et al., 2024) is a recently introduced GNN model that aims to learn and predict robust and transferable representations for positional and structural encodings using RNFs as inputs, and constitutes the main architecture studied in our paper. In this paper, we choose to work with GPSE because it generalizes other PSEs, such as the graph Laplacian eigenvectors or random-walk structural encodings, in the sense that it learns these PSEs during its training phase. In what follows, we provide a brief standalone introduction to the model; further information is available via the original paper.
GPSE model. GPSE employs an encoder-decoder architecture in which the encoder and decoder are trained jointly, whereas in inference, only the encoder is used. The encoder is a deep MPNN consisting of stacked GatedGCN (Bresson and Laurent, 2017) layers, with residual gating and skip-connections in-between layers. The decoder consists of multiple node-level MLP heads, each corresponding to a different PSE component.
with and are learnable parameters.
From here on we will assume that, given the structure of a graph , namely without any labels, as well as some random features , GPSE provides node embeddings Throughout most of this paper we consider GPSE to be pre-trained and thus fixed. We thus think of GPSE as a PSE that provides node embeddings.
PSEs. In this work, we will be comparing multiple PSEs, including GPSE. Here we give a brief overview of the most relevant PSEs. The Laplacian eigenvector positional encoding (LapPE) (Dwivedi et al., 2020; Lim et al., 2023) is the absolute value of the normalized eigenvector associated with non-trivial eigenvalues of the graph Laplacian. The random walk structural encoding (RWSE) (Dwivedi et al., 2022b) is the probability of returning back to the starting state of a random walk after steps. We refer to the combination of all PSEs used in Cantürk et al. (2024) as AllPSE. Lastly, we compare to Local-instance and Global-semantic Learning (GraphLog) (Xu et al., 2021), which constructs a locally smooth latent space by aligning the embeddings of correlated graphs/subgraphs. More details and some additional less relevant PSEs can be found in Appendix D,
Downstream universality. Throughout this work, we will be considering the expressivity that node augmentations (or positional and structural encodings) provide to GNNs. For example, as previously shown in Sato et al. (2021); Franks et al. (2023), random node features (RNF) added to GNN based on GIN layers enable the resulting model’s universality. In this sense, RNF itself is not universal but we will refer to this as “RNF enables downstream universality”. More generally, when we refer to the downstream universality or downstream expressivity of some PSE, then we are referring to the universality or expressivity that a GNN, that is able to approximate 1-WL, attains when its input features are augmented by this PSE.
@@ 4.1 GPSE is a Universal Node Encoder
Since GPSE is meant to be able to compute PSEs for any input graph and thus provide a powerful positional and structural node encoding, it is important to verify that GPSE is indeed a universal node encoder. Because some PSEs are harder—in a WL sense—to compute than CR, GPSE as an MPNN needs to be fully expressive. And indeed, GPSE is a universal node encoder, because the input graph is augmented with RNF, which we now show. First, note that GatedGCN layers can approximate GIN layers by the following theorem, which we prove in Appendix B.
@@
Combining Theorem 1 with Theorem 3 from Xu et al. (2019), which states that an MPNN made up of GIN layers can learn to approximate 1-WL, proves that GatedGCN can also simulate 1-WL. Furthermore, Theorem 4 from Franks et al. (2023) shows that, assuming sufficiently many layers and sufficient width of the layers within GPSE, any function on the graph structure (that is, functions of the form ignoring node/edge features) can be computed. This proves that, indeed, under mild assumptions, GPSE is universal and can learn to embed any PSE.
@@ 4.2 PSEs do not Guarantee Downstream Universality
While GPSE can learn to embed any PSE, this is not sufficient to ensure the universality of the downstream model. In fact, PSEs in general do not provide universality of the downstream model. As shown in Cantürk et al. (2024), as well as in Table 1, a GIN downstream network augmented with GPSE is not able to accurately generalize on the EXP, CEXP, CSL, TRI, and TRIX datasets, that are often used as expressivity benchmarks (Abboud et al., 2021; Murphy et al., 2019). Let us assume that the pre-trained network in GPSE learned an identity mapping of the input random node features. Then, the resulting model is theoretically universal with high probability, as shown in Abboud et al. (2021) and Sato et al. (2021). This would imply that the issue lies in the pre-training of GPSE. That is, GPSE did not learn to encode PSEs which obtain high accuracy on some graphs that require high expressivity to distinguish.
GPSE+. We empirically verify that by adding relevant “hard” graphs, i.e., graphs that require high expressivity, to the pre-training dataset of GPSE, GPSE is practically more expressive. Specifically, relevant hard graphs for the CSL dataset are 4-regular graphs. Thus we define GPSE+to be the exact same model as GPSE trained on MolPCBA (>300.000 unique graphs), as well as 1000 random 4-regular graphs, each with 24 nodes. We chose 24 nodes, as this is roughly the average number of nodes for MolPCBA. Indeed, Table 1 shows that, by adding these hard graphs to the pre-training stage, the performance of GPSE on CSL increases significantly. The performance on EXP also increases; however, since the added graphs are not particularly relevant to the EXP dataset, the performance does not change as dramatically as in the case of CSL. We also evaluated on TRI and TRIX, which are 3-regular graphs with triangles as labels, where again GPSE+outperforms GPSE. We note, that the use of 4-regular graphs was merely a demonstration focusing on the CSL dataset. In practice, either a task-tailored or a more general set of hard graphs would need to be used in the pretraining stage.
@@
GPSE-. While the GPSE+variant considers strengthening the pre-training stage with hard graphs, our experimental results in Section 5 on real-world datasets do not read improved performance. Therefore, to address the question of the importance of expressivity in GPSE, we also replace the input RNF with a constant feature vector, effectively removing the universality offered by RNF. We call this variant GPSE-. It has the same model architecture as GPSE; however, instead of random normally distributed inputs, it uses vectors, i.e., . In practice, this implies that the GPSE-model is not capable of computing all PSEs for any input graph. As an example, GPSE-cannot compute cycle counts for the graphs in Fig. 1(a) and Fig. 1(b), since GPSE-cannot tell these two graphs apart. However, since molecules are almost trees (Ahn et al., 2022), GPSE-should be able to still learn the target PSEs for the pre-training dataset, MolPCBA, as well as for the other evaluation datasets like ZINC-12k. Notably, Table 1 shows that GPSE-does not excel on datasets that involve graphs harder than 1-WL. In particular, the performance of GPSE-is roughly in line with a constant model. On TRIX GPSE-outperforms GPSE as well as GPSE+, because these other variants are having trouble generalizing well to the larger 3-regular graphs of TRIX, and actually perform worse than a naive, constant predictor. The evaluations in Section 5 will demonstrate that, indeed, RNF is not necessary for good practical performance either. In practice, at least for current datasets, GPSE, GPSE+, and GPSE-perform comparably.
Practical implications. As will be demonstrated in the next section, the differences between GPSE, GPSE+, and GPSE-in practice are virtually non-existent. This is likely due to the fact that expressivity is rarely relevant in practice. However, if expressivity is determined to be irrelevant GPSE-should probably be preferred, as the use of RNF in GPSE could lead to unpredictable results. While on the other hand, if expressivity is important, then an approach closer to GPSE+should be taken, specifically focusing on pretraining using hard graphs that are relevant to the task or generally hard graphs.